# QuantX: A Framework for Hardware-Aware Quantization of Generative AI Workloads

### v1

@@ Abstract
We present QuantX: a tailored suite of recipes for LLM and VLM quantization. It is capable of quantizing down to 3-bit resolutions with minimal loss in performance. The quantization strategies in QuantX take into account hardware-specific constraints to achieve efficient dequantization during inference ensuring flexible trade-off between runtime speed, memory requirement and model accuracy. Our results demonstrate that QuantX achieves performance within 6% of the unquantized model for LlaVa-v1.6 quantized down to 3-bits for multiple end user tasks and outperforms recently published state-of-the-art quantization techniques. This manuscript provides insights into the LLM quantization process that motivated the range of recipes and options that are incorporated in QuantX.
@@ 2.4 Hardware-Aware Constraints
The type and speed of software kernels that enable optimal usage of hardware features also feeds in deciding the optimal quantization strategy. LLM decoding is known to be a memory-bound operation and as such packing and unpacking kernels play an important role in determining the run-time. Recent SOTA quantization techniques [LZF+25, TCS+24] have suggested the use of transformation operations before quantization which can reduce the loss due to quantization. Such transformations, however, have to be reversed during dequantization which increases the runtime and reduces the achieved tokens per second. If maximizing inference speed is the main target then such techniques may not be suitable. On the other hand, if accuracy in limited memory is desired then custom dequantization kernels enable that. This, again, motivates the various hardware-informed design choices made in QuantX.
@@ 3 Results
In this section, we present our results on the VLM LlaVa-v1.6 7B [LLWL23] model. We compare the unquantized model in FP16 resolution with the quantized versions using the lmms-eval platform [ZLZ+24] on the Coco caption, VQAv2 and MMMU benchmarks. All of these benchmarks capture the ability of the model to translate visual content to text, as well as its visual and textual reasoning capability. In Table 1, we present results for our proposed QuantX quantization framework and AWQ [LTT+24], which has gained a lot of traction in the industry and been incorporated into several open-source projects. It can be seen that QuantX outperforms AWQ in all metrics considered and performs within 6% of the unquantized FP16 model. Further, this improvement in performance is achieved while using a smaller number of bits per weight (BPW) due to the larger group size in QuantX. We will update this paper with results on the runtime comparison of these algorithms in the future.
@@ 4 Conclusion
In this paper, we presented insights into the LLM quantization process that motivated the various design choices made in QuantX. We demonstrated that QuantX can successfully quantize LlaVa-v1.6 to 3-bits while achieving performance within 6% of the unquantized FP16 model. Furthermore, we showed that QuantX outperformed a SOTA technique while achieving a smaller BPW ratio. We intend to progressively update this manuscript with more data and insights.

### v2

@@ Abstract
We present QuantX: a tailored suite of recipes for LLM and VLM quantization. It is capable of quantizing down to 3-bit resolutions with minimal loss in performance. The quantization strategies in QuantX take into account hardware-specific constraints to achieve efficient dequantization during inference ensuring flexible trade-off between runtime speed, memory requirement and model accuracy. Our results demonstrate that QuantX achieves performance within 6% of the unquantized model for LlaVa-v1.6 quantized down to 3-bits for multiple end user tasks and outperforms recently published state-of-the-art quantization techniques. We further integrate one particular technique from QuantX into the popular Llama.cpp framework and show its feasibility in terms of runtime compared to the mainstream quantization techniques from Llama.cpp. Lastly, this manuscript provides insights into the LLM quantization process that motivated the range of recipes and options that are incorporated in QuantX.
@@ 2.4 Hardware-Aware Constraints
The type and speed of software kernels that enable optimal usage of hardware features also feeds in deciding the optimal quantization strategy. LLM decoding is known to be a memory-bound operation and as such packing and unpacking kernels play an important role in determining the run-time. Recent SOTA quantization techniques [LZF+25, TCS+24] have suggested the use of transformation operations before quantization which can reduce the loss due to quantization. Such transformations, however, have to be reversed during dequantization, which increases the runtime and reduces the achieved tokens per second. If maximizing inference speed is the main target, then such techniques may not be suitable. On the other hand, if accuracy in limited memory is desired, then custom dequantization kernels enable that. This, again, motivates the various hardware-informed design choices made in QuantX.
@@ 3 Results
In this section, we present our performance and run-time results on the VLM LlaVa-v1.6 7B [LLWL23] model and the LLM Llama-3.2-Instruct 1B [G+24].
3.1 Performance results
We first look at the VLM LlaVa-v1.6 7B [LLWL23] model and compare the unquantized model in FP16 resolution with the quantized versions using the lmms-eval platform [ZLZ+24] on the Coco caption, VQAv2 and MMMU benchmarks. All of these benchmarks capture the ability of the model to translate visual content to text, as well as its visual and textual reasoning capability. In Table 1, we present results for our proposed QuantX quantization framework and AWQ [LTT+24], which has gained a lot of traction in the industry and has been incorporated into several open source projects. It can be seen that QuantX outperforms AWQ in all metrics considered and performs within 6% of the unquantized FP16 model. Further, this improvement in performance is achieved while using a smaller number of bits per weight (BPW) due to the larger group size in QuantX. We will update this paper with results on the run-time comparison of these algorithms in the future.
@@ 3.1 Performance results
3.2 Runtime results
(a) Functional illustration of the Q4X quantization strategy showing the histogram learning process and non-uniform quantization for dimension reduction.
(b) Illustration of the quantized weight matrices for a single group of Q40, Q4K and Q4X.
Figure 3: (a) Description of the proposed Q4X quantization strategy and (b) the resulting quantized matrices for Q40, Q4K and Q4X.
Next, we demonstrate the practical feasibility of techniques in QuantX by integrating it into the popular Llama.cpp framework [GtLc25] and comparing it with the mainstream Llama.cpp techniques. Toward that end, we first give a brief overview of the built-in quantization strategies ‘Q40’ and ‘Q4K’ from Llama.cpp, followed by a breakdown of the new ‘Q4X’ quantization strategy from QuantX.
  • Q40: Q40 (shown on the left side of Fig. 3(b)) is a legacy 4-bit quantization method in Llama.cpp. It works by using a rounding-to-nearest (RTN) operation on a group of 32 weights. For each group, the 4-bit weights and an FP16 scaling factor are packed into 18 bytes achieving 4.5 BPW.
  • Q4K: Q4K is a more recent technique introduced in Llama.cpp that uses hierarchical scaling and asymmetric quantization. Here, 256 weight elements are grouped into a 144-byte block. This block contains 256 4-bit quantized values, a 2-byte global scale and minimum, and 6-bit scales and minimums for each chunk of 32 weights. Q4K also achieves a 4.5 BPW as illustrated in the middle of Fig. 3(b).
  • Q4X: As shown in Fig. 3(a), Q4X is a codebook-based quantization technique designed for efficient inference on CPUs. A histogram is computed for each group of 64 weights in the weight matrix followed by learning 4 histograms that best represent all the historgrams of a complete matrix. These learned histograms undergo dimension reduction, resulting in 16 non-uniform samples chosen from each. The final representation is made up of 64 centroids—16 from each of the 4 learned histograms—which are grouped into a codebook for the weight matrix. For each group of 64 weight elements, Q4X stores a 2-byte scale factor, a 2-bit codebook index that represents the best matching learned histogram for that group, and the 4-bit centroid indices for each of the 64 elements. This approach makes Q4X a 4.28 BPW (with the 2-bit codebook indices of multiple groups padded together for byte alignment) method, as illustrated on the right side of Fig. 3(b).
The final file size, run time performance in terms of tokens per second and perplexity computed on Wikitext v2 test subset with $n_{\text{ctx}}=2048$ is shown in Table 2. Q4X outperforms Q40, achieving a higher token throughput for a reduced model size and lower perplexity score. Similarly, Q4X has a better model size and throughput than Q4K. The perplexity score, however, is slightly lower for Q4K given its use of two-level scaling for the weights and a smaller sub-group size of 32 compared to the group size of 64 used in Q4X.
Quantization configuration BPW
File size
Prefill/Decode
throughput
(tokens/sec)
Perplexity
(Wikitext v2)
Unquantized FP16 16 $\sim$ 2300 0.12 / 0.06 13.16
Q40 W4A8G32 4.5 727.75 6.03 / 4.28 15.22
Q4K W4A8G32 4.5 732.25 7.77 / 5.06 14.45
Q4X W4A8G64 4.28 702.38 8.64 / 5.29 14.67
Table 2: Comparison of the file size, run time and perplexity for Llama 3.2 1B instruction tuned model and its Q40, Q4K and Q4X quantized versions running on a Milk-V Jupiter board. For a fair comparison in perplexity, we used 8-bit RTN per-token quantization for the activation tensors for all quantization techniques emulated in QuantX.
@@ 4 Conclusion
In this paper, we presented insights into the LLM quantization process that motivated the various design choices made in QuantX. We demonstrated that QuantX can successfully quantize LlaVa-v1.6 to 3-bits while achieving performance within 6% of the unquantized FP16 model. We also showed that QuantX outperformed a SOTA technique while achieving a smaller BPW ratio. Furthermore, we showed that QuantX extends the Pareto frontier across runtime, model size, and performance by integrating its 4-bit quantization into the Llama.cpp framework. We intend to progressively update this manuscript with more data and insights.